OpenStudy (anonymous):

Simplify \[\frac{sqrt{15}}{5}\] Please show all working

OpenStudy (anonymous):

\[\frac{\sqrt{15}}{5}\]

OpenStudy (anonymous):

I want to express it as \[\sqrt{\frac{3}{5}}\] but I am not sure how to.

OpenStudy (anonymous):

Do you know how to do it @dpaInc

OpenStudy (lgbasallote):

well that wouldnt be simplifying...

mathslover (mathslover):

\[\sqrt{15}/5\] \[\sqrt{5*3}/5\] \[\sqrt{5}*\sqrt{3}/5\] \[\sqrt{5}*\sqrt{5}*\sqrt{3}/5*\sqrt{5}\] \[5\sqrt{3}/5\sqrt{5}\] \[\sqrt{3}/\sqrt{5}\] \[\sqrt{3/5}\]

mathslover (mathslover):

this is the process through which u can represent \[\sqrt{15}/5 = \sqrt{3/5}\]

OpenStudy (lgbasallote):

but maybe this can give you a clue ;) \[\sqrt{\frac{3}{5}} \implies \frac{\sqrt 3}{\sqrt 5}\] now i multiply numerator and denominator by \(\sqrt 5\) \[\frac{\sqrt 3}{\sqrt 5} \times \frac{\sqrt 5}{\sqrt 5} = \frac{\sqrt{3 \times 5}}{\sqrt{5\times 5}} \implies \frac{\sqrt{15}}{5}\]

mathslover (mathslover):

hope u got it @purplec16

OpenStudy (callisto):

Perhaps this: Using the fact that\(\sqrt{25} =5\) \[\frac{\sqrt{15}}{5} =\frac{\sqrt{15}}{\sqrt{25}} = \sqrt{\frac{15}{25}} = \sqrt{\frac{5\times 3}{5\times 5}} = \sqrt{\frac{3}{5}}\]

OpenStudy (anonymous):

Yes, I did Thanks so much everyone!!!